In your own words define each of the following terms and give and short example:

1.Full capacity sales

2. Target fixed assets to sales ratio

3. Required level of fixed assets
Answer #1

1. full capacity sales is the sales happens when a 100% utilisation of resources in the business

example if a hotel has 10 rooms ,8 rooms are occupied ,2 rooms are vacant , let the room rent be 1000 per day

actual sales is 8 *1000 = 8000

full capacity sales is 10*1000 = 10000

2.target fixed assets to sales ratio = totalfixedasset / total sales  

how much the company possess the fixed asset regarding the revenue

for example x ownes a company named xyz and the total fixed asset is 40000 and revenue is 100000 then total fixed asset to sales ratio is 40000/100000 =0.4

3 required level of fixed asset is asset held for a longer period for the generation of profit ,required level of fixed asset may be needed for the budjeted profit so asset should be maintained according to the required level

example : for a paint company the avarage required level of fixed asset will be differ from hotel industry.it varies according to the level of business.